# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ Abstract
We present the latest results from the Chicago Carnegie Hubble Program ( CCHP) to measure the Hubble constant using data from the James Webb Space Telescope (JWST). This program is based upon three independent methods: (1) Tip of the Red Giant Branch (TRGB) stars, (2) JAGB (J-Region Asymptotic Giant Branch) stars, and (3) Cepheids. Our program includes 10 nearby galaxies, each hosting Type Ia supernovae (SNe Ia), suitable for measuring the Hubble constant (). It also includes NGC 4258, which has a geometric distance, setting the zero point for all three methods. The JWST observations have significantly higher signal-to-noise and finer angular resolution than previous observations with the Hubble Space Telescope (HST). We find three independent values of = 69.85 1.75 (stat) 1.54 (sys) for the TRGB, = 67.96 1.85 (stat) 1.90 (sys) km s-1 Mpc-1 for the JAGB, and = 72.05 1.86 (stat) 3.10 (sys) for Cepheids. Tying into SNe Ia, and combining these methods adopting a flat prior, yields our current estimate of = 69.96 1.05 (stat) 1.12 (sys) km s-1 Mpc-1. The distances measured using the TRGB and the JAGB method agree at the 1% level, but differ from the Cepheid distances at the 2.5-4% level. The value of based on these two methods with JWST data alone is = 69.03 1.75 (total error) km s-1 Mpc-1. These numbers are consistent with the current standard CDM model, without the need for the inclusion of additional new physics. Future JWST data will be required to increase the precision and accuracy of the local distance scale.
@@ 1 Introduction
The year 2024 marks just over a century since Edwin Hubble’s famous discovery of a single Cepheid variable in the Andromeda galaxy222The discovery plate was obtained 6 October 1923, on which Hubble wrote VAR! in red ink. See https://obs.carnegiescience.edu/PAST/m31var.. Hubble’s subsequent measurements of extragalactic distances (Hubble, 1929) were based, in part, on the Cepheid period-luminosity (PL) relation (also now widely known as the Leavitt Law, Leavitt (1908)). For 70 years a number of unrecognized challenges (e.g., reddening and dimming due to the effects of interstellar dust, errors in photometric zero points, effects due to differing metal abundances, crowding/blending as a result of insufficient resolution, and the inclusion of some (secondary) distance indicators) turned out to have large systematic effects, and combined together to make it virtually impossible for the Hubble consant () to be measured from the ground to better than a factor of two uncertainty.
@@
This impasse was largely overcome by technological advances: (1) the widespread availability of two-dimensional, linear detectors operating in the optical and near-infrared, beginning in the 1980s (e.g. McGonegal et al., 1982; Freedman et al., 1985, 1991), and (2) ultimately the launch of the Hubble Space Telescope (HST), and undertaking of the Hubble Key Project (Freedman et al.). These order-of-magnitude advances along the three orthogonal axes of sensitivity, wavelength coverage and angular resolution made it possible to reduce the 100% (‘factor-of-two’) uncertainty on the Hubble constant down to 10%, yielding = 72 3 (stat) 7 [sys] (Freedman et al., 2001). Over the following decades this space-based measurement of has been confirmed by multiple subsequent analyses (e.g., Riess et al., 2009; Freedman et al., 2012; Riess et al., 2016, 2022), all based on a Cepheid calibration of distant SNe Ia, subsequently taken into the more distant Hubble flow.
@@
Twenty-five years later, one of the outstanding questions in cosmology is this: Have we overcome the systematic effects (both known and perhaps still unknown) that can affect measurements of the astrophysical distance scale at a sufficient level to require that our current standard cosmological model (CDM) is now in need of additional physics? This question has arisen with the emergence of a completely new method for inferring the value of that is totally independent of the Cepheid calibration, and is, instead, based upon modeling measurements of the fluctuations in the cosmic microwave background. Under the assumption of the standard (CDM) cosmological model, the CMB measurements from the Planck satellite predict a current expansion rate of 67.4 0.5 km s-1 Mpc-1 (i.e., with better than 1% precision)(Planck Collaboration et al., 2020). Consistent results are also obtained from the Atacama Cosmology Telescope (ACT, (e.g., Madhavacheril et al., 2024)), and from the South Pole Telescope (SPT, (e.g., Balkenhol et al., 2023)). However, the CMB results differ from the most recent (local) measurements of based on distant Type Ia supernovae (SNe Ia) in the unperturbed Hubble flow, calibrated using HST observations of Cepheids. The difference is at a level of 5, as estimated by (Riess et al., 2022, hereafter R22), a discrepancy known as the Hubble tension. If this level of discrepancy is confirmed, it suggests the existence of new physics (particles or fields) not yet constrained by the standard (CDM) cosmological model. To date, however, no plausible changes or additions have emerged that allow for values of as high as 73 km s-1 Mpc-1 (see Di Valentino et al., 2021).
@@
Consistent with CMB measurements, lower values of are also attained from measurements of baryon acoustic oscillations (BAO), or fluctuations in the matter density. Most recently the Dark Energy Spectroscopic Instrument (DESI) Collaboration Data Release 1 (DR1) (DESI Collaboration et al., 2024) find = 68.52 0.62 km s-1 Mpc-1. This value was obtained by calibrating the standard BAO ruler based on a prior for the baryon density from Big Bang Nucleosynthesis (BBN), in addition to a measurement of the CMB acoustic angular scale, . (The CMB acoustic angular scale has been measured to a very high precision of 0.03% (Planck Collaboration et al., 2020)). A slightly less precise (but nonetheless the same) result is found using the BBN calibration alone, giving = 68.53 0.80 km s-1 Mpc-1. This result is completely independent of the CMB measurements. If instead, a calibration based on CMB measurement of the sound horizon distance, rd is used, the resulting value of is 69.29 0.87 km s-1 Mpc-1.
@@
HST has now served for more than 30 years as the primary, space-based instrument for the discovery and measurement of extragalactic Cepheid variables, and the local determination of (e.g., the Hubble Key Project (Freedman et al., 2001; Riess et al., 2022)). With its high resolution and sensitivity at optical wavelengths (where the amplitudes of Cepheids are large), HST is an ideal instrument for discovery of extragalactic Cepheids. In contrast, JWST is the best-equipped telescope for improving the accuracy of measurements of , due to its greater sensitivity at long wavelengths and its even higher spatial resolution relative to HST. The long-wavelength capability of JWST also makes it an ideal facility for the study of other distance indicators; for example, the redder TRGB and JAGB (J-Region Asymptotic Giant Branch) populations. Additionally, in the near- and mid-infrared, interstellar extinction is significantly lower than in the optical (e.g., AJ and A[4.4] are smaller by factors of 4 and 20 respectively, relative to the visual extinction, AV; and factors of 2 and 10 lower relative to the I-band extinction AI (Cardelli et al., 1989; Indebetouw et al., 2005). Moreover, the science performance of JWST has exceeded initial estimates of its sensitivity, stability, image quality, as well as spectral range (Rigby et al., 2022). NIRCam (F115W) imaging from JWST (Rieke et al., 2023) has a sampling resolution four times higher than HST (F160W), with a FWHM of 0.04 arcsec on JWST, versus 0.15 arcsec on HST. Of some potential concern, at red wavelengths, red giant and bright asymptotic giant branch stars can impact the photometry of the Cepheids (due to crowding and blending effects), exacerbating these effects in the red, compared to optical wavelengths. Importantly, with four times better resolution than HST at these near-infrared wavelengths, crowding effects are decreased by more than an order of magnitude in flux when using JWST.
@@
In this paper, we present our results from a new long-term Chicago-Carnegie Hubble Program (CCHP) using the James Webb Space Telescope (JWST). The aim of the program is to reduce the systematic uncertainty in the local extragalactic distance scale, and ultimately the measurement of . Our current goals are: (1) to use three independent stellar distance indicators (Cepheids, the TRGB, and JAGB stars) to obtain three high-precision distances to each calibrating galaxy, thereby reducing the overall systematic distance uncertainties, (2) to make use of the high resolution of JWST to understand and reduce the possible effects of crowding and blending of Cepheids in previous HST photometry, (3) to improve the corrections for dust, and (4) to improve constraints on the effects of metallicity on the Cepheid Leavitt law. Preliminary results from this program have been published in (Lee et al., 2024a; Owens+, 2024a, submitted; Hoyt et al., 2024; Freedman & Madore, 2023a, b). Details of the measurements for each of the methods discussed here are presented in three companion papers: (Owens+, 2024b, in prep; Hoyt+, 2024b, submitted; Lee+, 2024, submitted, hereafter O24, H24, L24, respectively). A fourth paper describes the details of the photometric reductions (Jang+, 2024, in prep). We note also that recent JWST results have also been reported by Yuan et al. (2022); Riess et al. (2023, 2024).
@@
The outline of this paper is as follows: In §2 we provide an overview of the CCHP. We then describe the galaxy sample and analysis in §3, our blinding procedure in §4, and the calibration in §5. In §6 we provide a description of the three methods used in this paper: Cepheids, the TRGB and the JAGB. In §7 we summarize our steps for the measurement of our new distances. In §8 we compare the distances obtained for the three methods, followed by a description of the calibration of SNe Ia and determination of in §9. In §10 we discuss the uncertainty that results from the small number of SN Ia calibrating galaxies, and §11 presents a discussion of the overall uncertainties and an error budget. In §12 we compare our results with previously published data. Section 13 provides a description of future prospects. Finally, in §14 we present our conclusions.
2 The Chicago Carnegie Hubble Program: An Overview
@@
In its second phase, the CCHP undertook a calibration of the Tip of the Red Giant Branch (TRGB) using the HSTAdvanced Camera for Surveys (ACS) to derive distances to nearby galaxies in which Type Ia supernovae (SNe Ia) had been discovered, and thus provided a calibration independent from that based on Cepheids alone. TRGB distances were measured to 15 galaxies that were known hosts to 18 SNe Ia. The halos of these galaxies were intentionally targeted to minimize the effects of dust, as well as to minimize contamination by younger and brighter disk asymptotic giant branch (AGB) stars. At the same time, crowding was also reduced by pre-selecting low surface brightness regions in the outer halo. The resulting distances were then used to tie into a more distant sample of SNe Ia, observed with high cadence and 9 filters (from the ultraviolet to the near infrared) as part of the Carnegie Supernova Project ( CSP) Krisciunas et al. (2017). This effort resulted in a value of = 69.8 0.6 (stat) 1.6 [sys] km s-1 Mpc-1 Freedman et al. (2019, 2020); Freedman (2021). The results were little changed if instead the SNe Ia catalog from the SHoES collaboration Scolnic et al. (2015) was adopted, giving = 70.4 1.4 1.6 km s-1 Mpc-1 Freedman et al. (2019).
In the current and third phase of the CCHP, we have undertaken a three-pronged approach to measuring using Cycle 1 JWST observations (JWST-GO-1995: P.I. W. L. Freedman). We observed 10 nearby galaxies that are hosts to type SNe Ia, and measured three independent distances using Cepheids, the TRGB and JAGB stars. In addition, we obtained observations of NGC 4258, a galaxy that is host to H2O megamasers, and which provides a geometric distance and our absolute calibration. The program was designed to deal specifically with known systematic effects in the measurement of distances to nearby galaxies: (a) extinction and reddening by dust along the total line of sight to these objects, (b) metallicity effects, and (c) crowding/blending of stellar images. Simply getting more nearby galaxy distances (decreasing the statistical uncertainties) is insufficient to confirm or refute whether new physics beyond the standard cosmological model is required. At this time, systematic uncertainties are (and have historically been) the dominant component of the error budget.

### v2

# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Measurement of the Hubble Constant Using the Hubble and James Webb Space Telescopes

@@ Abstract
We present the latest results from the Chicago-Carnegie Hubble Program ( CCHP) to measure the Hubble constant, using data from the James Webb Space Telescope (JWST). The overall program aims to calibrate three independent methods: (1) Tip of the Red Giant Branch (TRGB) stars, (2) JAGB (J-Region Asymptotic Giant Branch) stars, and (3) Cepheids. To date, our program includes 10 nearby galaxies, hosting 11 Type Ia supernovae (SNe Ia) suitable for measuring the Hubble constant (). It also includes the galaxy NGC 4258, whose geometric distance provides the zero-point calibration. In this paper we discuss our results from the TRGB and JAGB methods. Our current best (highest precision) estimate is = 70.39 1.22 (stat) 1.33 (sys) 0.70 (, based on the TRGB method alone, with a total of 24 SN Ia calibrators from both HST and JWST data. Based on our new JWST data only, and tying into SNe Ia, we find values of = 68.81 1.79 (stat) 1.32 (sys) for the TRGB, and = 67.80 2.17 (stat) 1.64 (sys) km s-1 Mpc-1 for the JAGB method. The distances measured using the TRGB and the JAGB method agree, on average, at a level better than 1%, and with the SHoES Cepheid distances at just over the 1% level. Our results are consistent with the current standard CDM model, without the need for the inclusion of additional new physics. Future JWST data will be required to increase the precision and accuracy of the local distance scale.
@@ 1 Introduction
The year 2025 marks a century since Edwin Hubble’s first publication of the discovery of Cepheid variables in an external galaxy, NGC 6822 (Hubble, 1925). Hubble’s subsequent measurements of extragalactic distances (Hubble, 1929) were based, in part, on the Cepheid period-luminosity (PL) relation (also now widely known as the Leavitt Law, Leavitt (1907)). For 70 years a number of unrecognized challenges (e.g., reddening and extinction/dimming due to the effects of interstellar dust, errors in photometric zero points, effects due to differing metal abundances, crowding/blending as a result of insufficient resolution, and the inclusion of some (secondary) distance indicators) turned out to have large systematic effects, and combined together made it virtually impossible for the Hubble constant () to be measured from the ground to better than a factor of two uncertainty.
@@
This impasse was largely overcome by technological advances: (1) the widespread availability of two-dimensional, linear detectors operating in the optical and near-infrared, beginning in the 1980s (e.g. McGonegal et al., 1982; Freedman et al., 1985, 1991), and (2) ultimately the launch of the Hubble Space Telescope (HST), and undertaking of the Hubble Key Project (Freedman et al., 2001). These order-of-magnitude advances along three axes of sensitivity, wavelength coverage and angular resolution, made it possible to reduce the 100% (‘factor-of-two’) uncertainty on the Hubble constant down to 10%, yielding = 72 3 (stat) 7 [sys] (Freedman et al.). Over the following decades this space-based measurement of has been confirmed by multiple subsequent analyses (e.g., Riess et al., 2009; Freedman et al., 2012; Riess et al., 2016, 2022), all based on a Cepheid calibration of distant SNe Ia, subsequently taken into the more distant Hubble flow.
@@
Twenty-five years later, one of the outstanding questions in cosmology is this: Have we overcome the systematic effects (both known and perhaps still unknown) that can affect measurements of the astrophysical distance scale at a sufficient level to require that our current standard cosmological model (CDM) might now be in need of additional physics? This question has arisen with the emergence of a completely new method for inferring the value of that is totally independent of the Cepheid calibration, and is, instead, based upon modeling measurements of the fluctuations in the cosmic microwave background. Under the assumption of the standard (CDM) cosmological model, the CMB measurements from the Planck satellite predict a current expansion rate of 67.4 0.5 km s-1 Mpc-1 (i.e., with better than 1% precision)(Planck Collaboration et al., 2020). Consistent results are also obtained from the Atacama Cosmology Telescope (ACT, (e.g., Madhavacheril et al., 2024)), and from the South Pole Telescope (SPT, (e.g., Balkenhol et al., 2023; Ge et al., 2024)). However, the CMB results differ from some recent (local) measurements of based on distant Type Ia supernovae (SNe Ia) in the unperturbed Hubble flow, calibrated using HST observations of Cepheids. The difference is at a level of 5, as estimated by (Riess et al., 2022, hereafter R22), a discrepancy known as the Hubble tension. However, HST TRGB measurements (Freedman et al., 2019) show no significant tension with the CMB. If a large discrepancy is confirmed, it suggests the existence of new physics (particles or fields) not yet constrained by the standard (CDM) cosmological model. To date, however, no plausible changes or additions have emerged that allow for values of as high as 73 km s-1 Mpc-1 (see Di Valentino et al., 2021).
@@
Other physics-based methods, consistent with CMB measurements and lower values of , are also attained from measurements of baryon acoustic oscillations (BAO), or fluctuations in the matter density. The Dark Energy Spectroscopic Instrument (DESI) Collaboration Data Release 1 (DR1) (DESI Collaboration et al., 2024) find = 68.52 0.62 km s-1 Mpc-1. This value was obtained by calibrating the standard BAO ruler based on a prior for the baryon density from Big Bang Nucleosynthesis (BBN), in addition to a measurement of the CMB acoustic angular scale, . (The CMB acoustic angular scale has been measured to a very high precision of 0.03% (Planck Collaboration et al., 2020)). A slightly less precise (but nonetheless the same) result is found using the BBN calibration alone, giving = 68.53 0.80 km s-1 Mpc-1. While not completely independent of the CMB measurements (these results assume the CDM model), they offer an important consistency check. If instead, a calibration based on CMB measurement of the sound horizon, rd is used, the resulting value of is 69.29 0.87 km s-1 Mpc-1 (again assuming the CDM model). Most recently, and also based on the DESI data, combined with additional sound-horizon independent data sets, Zaborowski et al. (2024) find km s-1 Mpc-1.
Although modeling of gravitationally lensed supernovae is in its early stages, measurements of the well-modeled) Supernova Refsdal have yielded low values of = 64.8 4.4 and 66.6 4.4 km s-1 Mpc-1 (Kelly et al., 2023); = 65.1 km s-1 Mpc-1 (Grillo et al., 2024)222A second gravitational lens SN H0pe (Pascale et al., 2025) currently has almost twice the uncertainty 9%.. These several, independent (physics-based) methods (CMB, BAO, lensed supernovae) provide no evidence for new early-Universe physics.
Beyond the challenges of calibrating an absolute distance scale, accurate local measurements based on SNe Ia necessitate precise calibration of SN Ia peak absolute magnitudes (MB). A well-established degeneracy exists between MB and H0 (e.g., Efstathiou, 2021; Dainotti et al., 2025). In essence, any error in MB whether stemming from measurements of nearby SNe Ia or the distances to Cepheids or TRGB stars used for calibration, directly propagates into an error in the measured value. Literature values for MB range from -19.253 0.027 (Riess et al., 2022) calibrated by Cepheids, to -19.396 0.016 (Dinda & Banerjee, 2023), based on a combination of SNe Ia and BAO observations. A calibration of MB using the TRGB yields MB = -19.326 0.038 mag (Freedman et al., 2019) for a sample of galaxies overlapping the Cepheid sample of Riess et al. (2016). Chen et al. (2024) have analyzed the Pantheon+ SNe Ia sample that forms the basis of the Riess et al. (2022) study, and illustrate that the published range of choices for values of MB result in variations in the mean values ranging from 2% to 7% (i.e., up to 5 km s-1 Mpc-1.) A critical issue remains the small sample size of nearby SNe Ia available for absolute calibration in the extragalactic distance scale (Riess et al., 2022). Furthermore, different groups employing varying SN Ia light-curve fitting techniques have yet to achieve convergence in their luminosity measurements (e.g., Vincenzi et al., 2025, and references therein). Resolving the absolute magnitude calibration for SNe Ia is essential for achieving high-accuracy measurements.
@@
HST has now served for more than 30 years as the primary, space-based instrument for the discovery and measurement of extragalactic Cepheid variables, and the local determination of (e.g., the Hubble Key Project (Freedman et al., 2001), and SHoES (Riess et al., 2022)). With its high resolution and sensitivity at optical wavelengths (where the amplitudes of Cepheids are large), HST is an ideal instrument for discovery of extragalactic Cepheids. In contrast, JWST is the best-equipped telescope for improving the accuracy of measurements of , due to its greater sensitivity at long wavelengths (where Cepheid amplitudes are generally lower) and its even higher spatial resolution relative to HST. The long-wavelength capability of JWST also makes it an ideal facility for the study of other distance indicators; for example, the redder TRGB and JAGB (J-Region Asymptotic Giant Branch) populations. Additionally, in the near- and mid-infrared, interstellar extinction is significantly lower than in the optical (e.g., AJ and A[4.4] are smaller by factors of 4 and 20, respectively, relative to the visual extinction, AV; and factors of 2 and 10 lower relative to the I-band extinction, AI (Cardelli et al., 1989; Indebetouw et al., 2005). Moreover, the science performance of JWST has exceeded initial estimates of its sensitivity, stability, image quality, as well as spectral range (Rigby et al., 2022). NIRCam (F115W) imaging from JWST (Rieke et al., 2023) has a sampling resolution four times higher than HST (F160W), with a FWHM of 0.04 arcsec on JWST, versus 0.15 arcsec on HST. Of some potential concern, at red wavelengths, red giant and bright asymptotic giant branch stars can impact the photometry of the Cepheids (due to crowding and blending effects), exacerbating these effects in the red, compared to optical wavelengths. Importantly, with four times better resolution than HST at these near-infrared wavelengths, crowding effects are decreased by more than an order of magnitude in flux when using JWST.
@@
In this paper, we present our results from a new long-term Chicago-Carnegie Hubble Program (CCHP) using the James Webb Space Telescope (JWST). The aim of the program is to reduce the systematic uncertainties in the local extragalactic distance scale, and ultimately to provide a robust measurement of . Our goals are: (1) to use three independent stellar distance indicators (Cepheids, the TRGB, and JAGB stars) to obtain three high-precision JWST-only distances to each calibrating galaxy, thereby reducing the overall systematic distance uncertainties, (2) to make use of the high resolution of JWST to understand and reduce the possible effects of crowding and blending of Cepheids in previous HST photometry, (3) to improve the corrections for dust, and (4) to improve constraints on the effects of metallicity on the Cepheid Leavitt law. Preliminary results from this program have been published in (Lee et al., 2024a; Owens et al., 2025a; Hoyt et al., 2024; Freedman & Madore, 2023a, b). Details of the measurements for the TRGB and JAGB results discussed in this paper are presented in two companion papers: Lee et al. (2024b), Hoyt et al. (2025) [hereafter L25 and H25, respectively]. A third paper on the Cepheid distance scale (Owens et al., 2025b) is in preparation. We note also that recent JWST results have also been reported by Yuan et al. (2022); Riess et al. (2023); Li et al. (2024); Riess et al. (2024a, b); Li et al. (2025).
@@
The outline of this paper is as follows: In §2 we provide an overview of the CCHP. We then describe the galaxy sample and analysis in §3, our blinding procedure in §4, and our calibration in §5. In §6 we provide a description of the two methods used in this paper: the TRGB and the JAGB. In §7 we summarize our steps for the measurement of our new distances. In §8 we compare the distances obtained for the two methods, followed by a comparison of previous TRGB measurements in §9. A description of the calibration of SNe Ia and determination of is given in §10. In §11 we discuss the uncertainty that results from the small number of SN Ia calibrating galaxies. We discuss the current status of the Hubble tension in §12. Section 13 presents a discussion of the overall uncertainties and an error budget. In §14 we compare our results with previously published data. Section 15 provides a description of future prospects. Finally, in §16 we present our summary and conclusions. We anticipate the main results of this study in Figure 1.
2 The Chicago-Carnegie Hubble Program: An Overview
@@
In its second phase, the CCHP undertook a calibration of the Tip of the Red Giant Branch (TRGB) using the HSTAdvanced Camera for Surveys (ACS) to derive distances to nearby galaxies in which Type Ia supernovae (SNe Ia) had been discovered, and thus provided a calibration that was independent from that based on Cepheids alone. TRGB distances were measured to 15 galaxies that were known hosts to 18 SNe Ia. The halos of these galaxies were intentionally targeted to minimize the effects of dust, as well as to minimize contamination by younger and brighter disk asymptotic giant branch (AGB) stars. At the same time, crowding was also reduced by pre-selecting low surface brightness regions in the outer halo. The resulting distances were then used to tie into a more distant sample of SNe Ia, observed with high cadence and 9 filters (from the ultraviolet to the near infrared) as part of the Carnegie Supernova Project ( CSP) Krisciunas et al. (2017). This effort resulted in a value of = 69.8 0.6 (stat) 1.6 (sys) km s-1 Mpc-1 (Freedman et al., 2019, 2020; Freedman, 2021). The results were little changed if instead the SNe Ia catalog from the SHoES collaboration Scolnic et al. (2015) was adopted, giving = 70.4 1.4 km s-1 Mpc-1 (Freedman et al., 2019).
In the current and third phase of the CCHP, we have undertaken a three-pronged approach to measuring using Cycle 1 JWST observations (JWST-GO-1995: P.I. W. L. Freedman). We observed 10 nearby galaxies that are hosts to 11 type SNe Ia, and measured three independent distances using Cepheids, the TRGB and JAGB stars. In addition, we obtained observations of NGC 4258, a galaxy that is host to H2O megamasers, and which provides a geometric distance and our absolute calibration. The program was designed to deal specifically with known systematic effects in the measurement of distances to nearby galaxies: (a) extinction and reddening by dust along the total line of sight to these objects, (b) metallicity effects, and (c) crowding/blending of stellar images. Simply getting more nearby galaxy distances (decreasing the statistical uncertainties) is insufficient to confirm or refute whether new physics beyond the standard cosmological model is required. At this time, systematic uncertainties are (and have historically always been) the dominant component of the error budget.